### v1

# Concentration of a sparse Bayesian model with Horseshoe prior in estimating high-dimensional precision matrix

Abstract
Precision matrices are crucial in many fields such as social networks, neuroscience, and economics, representing the edge structure of Gaussian graphical models (GGMs), where a zero in an off-diagonal position of the precision matrix indicates conditional independence between nodes. In high-dimensional settings where the dimension of the precision matrix $p$ exceeds the sample size $n$ and the matrix is sparse, methods like graphical Lasso, graphical SCAD, and CLIME are popular for estimating GGMs. While frequentist methods are well-studied, Bayesian approaches for (unstructured) sparse precision matrices are less explored. The graphical horseshoe estimate by Li et al., [27], applying the global-local horseshoe prior, shows superior empirical performance, but theoretical work for sparse precision matrix estimations using shrinkage priors is limited. This paper addresses these gaps by providing concentration results for the tempered posterior with the fully specified horseshoe prior in high-dimensional settings. Moreover, we also provide novel theoretical results for model misspecification, offering a general oracle inequality for the posterior.
62J12,
62F15,
62J99,
62C10,
Precision matrix,
high-dimensional data,
Posterior concentration rate,
sparsity,
misspecified model,
horseshoe prior,
keywords:
[class=MSC]
keywords:
[…]
1 Introduction
Precision matrices play a pivotal role in numerous fields, including social networks, neuroscience, and economics [33, 35, 16, 11]. The configuration of the precision matrix $\Omega=\Sigma^{-1}\in\mathbb{R}^{p\times p}$ represents the edge structure of a Gaussian graphical model, where the nodes are normally distributed as $\mathcal{N}(0,\Sigma)$ [25]. Gaussian graphical models (GGMs) continue to be favored for network estimation due to the straightforward interpretation of the precision matrix: a zero in an off-diagonal position indicates conditional independence between the two corresponding nodes, given all other nodes. Due to these important implications, the estimation of precision matrices has received substantial attention over the past few decades.
In practical applications, users often encounter situations where the dimension $p$ of the precision matrix is comparable to or even exceeds the sample size $n$, and the precision matrix exhibits sparsity. In such contexts, leveraging the sparsity of the precision matrix is crucial to discern the edge structure of the graph and assess the conditional dependencies among the nodes. Among the most widely used frequentist methods for estimating GGMs are the graphical lasso [17], the graphical SCAD [15], and the CLIME estimator [9]. For a comprehensive overview, see Cai et al., [10] and Janková and van de Geer, [23]. These methods provide estimates of high-dimensional inverse covariance matrices under various sparsity patterns.
The body of work on Bayesian methodologies for unstructured precision matrices is relatively limited. Wang, [40] introduced a Bayesian variant of the graphical lasso, utilizing block Gibbs sampling. Banerjee and Ghosal, [5] employed a similar prior structure to the Bayesian graphical lasso but included a substantial point-mass at zero for the off-diagonal elements of the precision matrix. They derived posterior convergence rates in the Frobenius norm under sparsity assumptions and proposed a Laplace approximation method for calculating marginal posterior probabilities of models. Additionally, Gan et al., [19] proposed spike-and-slab variants with double exponential priors. Atchadé, [3] combined the spike-and-slab prior with a pseudo-likelihood approach, resulting in a quasi-posterior distribution with favorable contraction properties.
A common issue with the spike-and-slab approach is the presence of binary indicator variables, which can hinder posterior exploration. The graphical horseshoe method, introduced by Li et al., [27], addresses these issues by applying the global-local horseshoe prior [12] to Gaussian graphical models. Li et al., [27] provided substantial empirical evidence of the graphical horseshoe estimate’s superior performance compared to several competing Bayesian and frequentist methods in both low and high-dimensional settings. However, there is limited research on the theoretical properties of shrinkage priors for sparse precision matrix estimation. Recent studies by Zhang et al., [43] and Sagar et al., [36] have adapted several versions of Horseshoe prior in different ways but do not provide theoretical justification for the method proposed by Li et al., [27]. Specifically, Zhang et al., [43] consider a pseudo-likelihood approach similar to Atchadé, [3], while Sagar et al., [36] explore a horseshoe-like prior. Moreover, both studies offer theoretical results only in an asymptotic sense and do not address high-dimensional settings. This work provides an attempt to address these gaps in the literature.
In this paper, we provide concentration results for the posterior in estimating precision matrices utilizing the horseshoe prior as in Li et al., [27]. Following Banerjee and Ghosal, [5], Zhang et al., [43], and Sagar et al., [36], we assume that the underlying precision matrix is sparse. Importantly, our results are derived in a high-dimensional setting where $p>n$. Moreover, while Zhang et al., [43] and Sagar et al., [36] consider either a horseshoe-like prior or a horseshoe prior with a fixed global shrinkage parameter, our results are obtained for the fully specified horseshoe prior used in practice as in Li et al., [27]. Another novel aspect of our work is that we also provide theoretical results for the case of model misspecification. Specifically, we prove a general oracle inequality for the posterior under an abstract setting that allows for model misspecification.
The rest of the paper comprises the following sections Section 2 present the problem of sparse precision matrix estimation, a Bayesian approach using tempered posterior and the Horseshoe prior. Section 3 contains our main results. We conclude in Section 4. All technical proofs re gathered in Appendix A.
2 Problem and method
Notations
Let $P,R$ be two probability measures and $\mu$ be any measure such that $P\ll\mu$ and $R\ll\mu$. For $\alpha\in(0,1)$, the $\alpha$-Rényi divergence between two probability distributions $P$ and $R$ is defined by
$D_{\alpha}(P,R)=\frac{1}{\alpha-1}\log\int\left(\frac{{\rm d}P}{{\rm d}\mu}\right)^{\alpha}\left(\frac{{\rm d}R}{{\rm d}\mu}\right)^{1-\alpha}{\rm d}\mu\text{,}$
and the Kullback-Leibler divergence is defined by
$\mathcal{K}(P,R)=\int\log\left(\frac{{\rm d}P}{{\rm d}R}\right){\rm d}P$
[…]
2.1 Model
We are interested in estimating the precision matrix $\Omega\in\mathbb{R}^{p\times p}$ from the sample data $Y\in\mathbb{R}^{n\times p}$, where $n$ denotes the sample size, and $p$ is the dimension of precision matrix or equivalently the number of nodes in the corresponding graph. The rows of $Y$ are assumed to be i.i.d. $\mathcal{N}(0,\Omega^{-1})$ distributed. The likelihood function based on the data can be written as
$L(Y|\Omega)=(2\pi)^{-\frac{np}{2}}\det(\Omega)^{\frac{n}{2}}\exp\left\{-\frac{1}{2}tr(Y\Omega Y^{T})\right\}.$ (2.1)
Similar to [3, 43, 36], we assume the diagonal elements $\omega_{ii}$ of $\Omega$ are known. Without loss of generality, we assume the diagonal $\omega_{ii}=1$ for all $i=1,\ldots,p$. In this work, we consider high-dimensional setting that $p>n$. We assume that the true precision matrix $\Omega_{0}$ is sparse in the sense that its number of non-zero off-diagonal elements is small.
2.2 Method
For a Bayesian model as in [27], the horseshoe prior [12] is independently specified on the elements of $\Omega$ as
$\omega_{ij}\mid\lambda_{ij},\tau\sim N(0,\lambda_{ij}^{2}\tau^{2}),\quad\lambda_{ij}\sim\mbox{Ca}_{+}(0,1),\quad\tau\sim\mbox{Ca}_{+}(0,1),$ (2.2)
for $1\leq i<j\leq p$, where $\mbox{Ca}_{+}(0,1)$ denotes the truncated standard half-Cauchy distribution with density proportional to $(1+t^{2})^{-1}\mathbbm{1}_{(0,\infty)}(t)$. We shall denote this prior induced by the hierarchy in (2.2) by $\Pi_{HS}$.
It is important to note that [43] presents theoretical results exclusively for the case where $\tau$ is a fixed value. However, in their numerical studies, they use a half-Cauchy prior, as in [27]. Therefore, our work extends their theoretical framework.
We focus on the tempered posterior given as
$\Pi_{n,\alpha}(\Omega)\propto L(Y|\Omega)^{\alpha}\Pi_{HS}(\Omega).$
for $\alpha\in(0,1)$. It is also called as fractional posterior [7, 2]. The regular posterior is obtained when $\alpha=1$, however, using a smaller value in general would be benifit in both computational [18] or in model mis-specified cases [21]. It is worth mentioning that tempered/fractional posteriors has received significant attention in recent years, as in [1, 8, 13, 38, 28, 29, 30, 31, 42, 41].
Put
$\hat{\Omega}=\int\Omega\Pi_{n,\alpha}({\rm d}\Omega)$
as our mean estimator.
3 Results
3.1 Concentration result for Horseshoe prior
The subsequent lemma introduces a novel prior concentration result for the horseshoe prior, offering a lower bound on the probability assigned to a Euclidean neighborhood of a sparse vector. This lemma represents a slight improvement over Lemma 1 in [13], crucial for achieving the optimal rate $s\log(p/s)/n$ instead of the sub-optimal rate $s\log(p)/n$ in high dimensional setting, see [6], [20].
Lemma 1.
Suppose $\theta_{0}\in\mathbb{R}^{d}$ such that $\|\theta_{0}\|_{0}=s$ and that $s<n<d$ and $\|\theta_{0}\|_{\infty}\leq C_{1}$. Suppose $\theta\sim\Pi_{HS}$. Define $\delta=\{s\log(d/s)/n\}^{1/2}$. Then we have, for some positive constant $K$,
$\Pi_{HS}(\|\theta-\theta_{0}\|_{2}<\delta)\geq e^{-Ks\log(d/s)}.$
This lemma play a key ingredient of our theory in high-dimensional sparse settings. Similar results but for different sparsity priors can found for example in [14, 32, 34].
3.2 Consistency of tempered posterior
Let $p_{\Omega}$ and $P_{\Omega}$ denote the density and the distribution of a Gaussian $\mathcal{N}(0,\Omega^{-1})$ random variable. Initially, we enumerate our assumptions regarding the truth.
Assumption 1 (Bounded spectrum).
The true precision matrix $\Omega_{0}$ satisfies that there exists a universal constant $\varepsilon_{0}>0$ that
$0<\varepsilon_{0}^{-1}\leq\mathrm{eig}_{\rm min}(\Omega)\leq\mathrm{eig}_{\rm max}(\Omega)\leq\varepsilon_{0}<\infty.$
Assumption 2.
The true precision matrix $\Omega_{0}$ is sparse in the sense that
$\Omega_{0}\in\mathcal{S}(s):=\left\{\Omega:\!\!\sum_{1\leq i<j\leq p}\mathbf{1}_{(\omega_{ij}\neq 0)}\leq s\right\}.$
Assumption 3 (Sparsity).
It is assumed that $s\log(p/s)/n=o(1)$.
Assumption 1 and 2 are standard in both frequentist and Bayesian literature [10, 23, 5, 36, 43]. Assumption 3 is a sparsity condition often assumed in high-dimensional settings [6, 20, 4]. We now present first a consistency result of the tempered posterior in $p>n$ setting.
Theorem 3.1.
For any $\alpha\in(0,1)$, assume that Assumptions 1, 2, and 3 are satisfied. Given any $\Omega_{0}$ such that $\|\Omega_{0}\|_{\infty}\leq C_{1}$, then
$\mathbb{E}\left[\int D_{\alpha}(P_{\Omega},P_{\Omega_{0}})\Pi_{n,\alpha}({\rm d}\Omega)\right]\leq\frac{1+\alpha}{1-\alpha}\varepsilon_{n},$
where $\varepsilon_{n}=C_{\varepsilon_{0}}\frac{s\log(p/s)}{n},$ and $C_{\varepsilon_{0}}$ is a universal positive constant depending only on $\varepsilon_{0}$.
Our proof technique leverages the general concentration theory for tempered posteriors as described in Bhattacharya et al., [7] and Alquier and Ridgway, [2], which ensures the desired convergence with respect to the $\alpha$-Rényi divergence. However, in the context of precision matrices, the Frobenius norm is more interpretable than the $\alpha$-Rényi divergence. The paper Banerjee and Ghosal, [5] demonstrated that, given appropriate assumptions on the eigenspace of precision matrices, the Frobenius norm and the Hellinger distance are equivalent. Additionally, Van Erven and Harremos, [39] established connections between the Hellinger distance and the $\alpha$-Rényi divergence. Therefore, the same results apply when using the Frobenius norm that is stated in the following corollary.
Put $c_{\alpha}:=1$ for $\alpha\in[0.5,1)$ and $c_{\alpha}:=(1-\alpha)/\alpha$ for $\alpha\in(0,0.5)$.
Corollary 3.1.
Assuming that Theorem 3.1 holds, then there exists a positive constant $c_{0}$ such that
$\mathbb{E}\left[\int\|\Omega-\Omega_{0}\|_{2}^{2}\Pi_{n,\alpha}({\rm d}\Omega)\right]\leq c_{0}c_{\alpha}\frac{1+\alpha}{1-\alpha}\varepsilon_{n},$
and an application of Jensen’s inequality,
$\mathbb{E}\left[\|\hat{\Omega}-\Omega_{0}\|_{2}^{2}\right]\leq c_{0}c_{\alpha}\frac{1+\alpha}{1-\alpha}\varepsilon_{n}.$
3.3 Concentration rates
We now provide our main result concerning the concentration of the fractional posterior relative to the $\alpha$-Rényi divergence of the densities.
Theorem 3.2.
Given any $\alpha\in(0,1)$, assume that Assumptions 1, 2, and 3 are satisfied. For all $\Omega_{0}$ such that $\|\Omega_{0}\|_{\infty}\leq C_{1}$, then
$\mathbb{P}\left[\int D_{\alpha}(P_{\Omega},P_{\Omega_{0}})\Pi_{n,\alpha}({\rm d}\Omega)\leq\frac{2(\alpha+1)}{1-\alpha}\varepsilon_{n}\right]\geq 1-\frac{2}{n\varepsilon_{n}},$
where $\varepsilon_{n}=C_{\varepsilon_{0}}\frac{s\log(p/s)}{n},$ and $C_{\varepsilon_{0}}$ is a universal positive constant depending only on $\varepsilon_{0}$.
We remind that all technical proofs are given in Appendix A.
Remark 1.
Our results imply that the concentration rate $\varepsilon_{n}$ are adaptive to the unknown sparsity level $s$ of the true precision matrix. The concentration rate is of order $s\log(p/s)/n$ is known to be minimax-optimal in high-dimensional setting [6, 20]. Readers may prefer [4] for a recent review of Bayesian methods in different high-dimensional models.
The same results apply when using the Frobenius norm that is stated in the following corollary.
Corollary 3.2.
As a special case, Theorem 3.2 leads to
$\mathbb{P}\left[\int\|\Omega-\Omega_{0}\|_{2}^{2}\,\Pi_{n,\alpha}({\rm d}\Omega)\leq 2c_{\alpha}\frac{1+\alpha}{1-\alpha}\varepsilon_{n}\right]\geq 1-\frac{2}{n\varepsilon_{n}},$
$\mathbb{P}\left[\|\hat{\Omega}-\Omega_{0}\|_{2}^{2}\leq 2c_{\alpha}\frac{1+\alpha}{1-\alpha}\varepsilon_{n}\right]\geq 1-\frac{2}{n\varepsilon_{n}}.$
Corollary 3.2 shows that the tempered posterior distribution of $\Omega$ concentrates around its true value, $\Omega_{0}$, at a specified rate relative to the squared Frobenius norm. As follow with previous works on tempered posterior [13, 2, 7], our results do not require that the true value can be tested against sufficiently separated other values in some suitable sieve as in [36] and [43].
One may note that our tempered posterior estimates are not necessarily positive definite. In some cases, a positive semi-definite (PSD) matrix may be more preferable. To obtain a PSD matrix $\tilde{\Omega}$, a post-processing step from [43] can be used:
$\tilde{\Omega}:=\operatorname*{arg\,min}_{\Omega\text{ is PSD }}\|\Omega-\bar{\Omega}\|_{2},$
where $\bar{\Omega}$ represents random samples from the posterior distribution $\Pi_{n,\alpha}(\Omega)$. As shown in [22], this optimization problem has a closed-form solution given by $\tilde{\Omega}=(B+H)/2$, where $H$ is the symmetric polar factor of $B=(\bar{\Omega}+\bar{\Omega}^{\top})/2$. This leads to the inequality $\|\tilde{\Omega}-\Omega_{0}\|_{2}\leq 2\|\Omega-\Omega_{0}\|_{2},$ which ensures the posterior concentration for the PSD matrix $\tilde{\Omega}$, stated in the following Corollary.
Corollary 3.3.
Assume that Theorem 3.2 holds, then
$\mathbb{P}\left[\int\|\tilde{\Omega}-\Omega_{0}\|_{2}^{2}\,\Pi_{n,\alpha}({\rm d}\Omega)\leq 2c_{\alpha}\frac{1+\alpha}{1-\alpha}\varepsilon_{n}\right]\geq 1-\frac{2}{n\varepsilon_{n}}.$
3.4 Result in the misspecified case
In this section, we show that our previous results can be extended to the misspecified setting. Assume that the true data generating distribution is parametrized by $\Omega_{0}\notin\mathcal{S}(s)$ and define $P_{\Omega_{0}}$ as the true distribution. Put
$\Omega_{*}:=\arg\min_{\Omega\in\mathcal{S}(s)}\mathcal{K}(P_{\Omega_{0}},P_{\Omega}),$
In order not to change all the notation we define an extended parameter set $\{\Omega_{0}\}\cup\mathcal{S}(s)$. Here, we clearly state now that that only $\Omega_{*}$ is in $\mathcal{S}(s)$. Example for this circumstance is that $\Omega_{0}$ may no longer be sparse.
Assumption 4.
The minimal precision matrix $\Omega_{*}$ satisfies that
[…]
Theorem 3.3.
For any $\alpha\in(0,1)$, let assume that Assumption 4 and 3 hold. For all $\Omega_{0}$ such that $\|\Omega_{*}\|_{\infty}\leq C_{1}$, then
$\mathbb{E}\left[\int D_{\alpha}(P_{\Omega},P_{\Omega_{0}})\Pi_{n,\alpha}({\rm d}\Omega)\right]\leq\frac{\alpha}{1-\alpha}\min_{\Omega\in\mathcal{S}(s,\varepsilon_{0})}\mathcal{K}(P_{\Omega_{0}},P_{\Omega})+\frac{1+\alpha}{1-\alpha}\varepsilon_{n},$
where $\varepsilon_{n}=C_{\varepsilon_{0}}\frac{s\log(p/s)}{n},$ and $C_{\varepsilon_{0}}$ is a universal positive constant depending only on $\varepsilon_{0}$.
In the scenario that $\Omega_{0}\in\mathcal{S}(s,\varepsilon_{0})$, i.e. the model is well-defined, meaning that $\mathcal{K}(P_{\Omega_{0}},P_{\Omega})=0$ over $\mathcal{S}(s,\varepsilon_{0})$, Theorem 3.1 is a special case of Theorem 3.3. Otherwise, the outcome provides a general oracle inequality.
Although it is not a sharp oracle inequality due to the differing risk measures on both sides, this observation remains valuable, particularly when $\mathcal{K}(P_{\theta_{0},\sigma_{0}},P_{\theta^{*},\sigma})$ is minimal. Nonetheless, under additional assumptions, we can further derive an oracle inequality result with $\ell_{2}$ distance on both sides. The result is as follows.
Corollary 3.4.
Assume that Theorem 3.3 holds . Then, there exist a universal constant $C>0$ such that
$\mathbb{E}\left[\int\|\Omega-\Omega_{0}\|_{2}^{2}\,\Pi_{n,\alpha}({\rm d}\Omega)\right]\leq\frac{\alpha}{1-\alpha}\min_{\Omega\in\mathcal{S}(s,\varepsilon_{0})}C\|\Omega-\Omega_{0}\|_{2}^{2}+\frac{1+\alpha}{1-\alpha}\varepsilon_{n},$
$\mathbb{E}\left[\|\hat{\Omega}-\Omega_{0}\|_{2}^{2}\right]\leq\frac{\alpha}{1-\alpha}\min_{\Omega\in\mathcal{S}(s,\varepsilon_{0})}C\|\Omega-\Omega_{0}\|_{2}^{2}+\frac{1+\alpha}{1-\alpha}\varepsilon_{n}.$
To the best of our knowledge, results in Corollary are completely novel for estimating sparse precision matrix from a Bayesian perspective.
The optimal constant may be achieved at $\alpha=1/2$. This aligns with the optimality of the half-power found in [26] within the framework of a pseudo-likelihood approach for model aggregation of least squares estimates, which also has a Bayesian interpretation as a fractional posterior.
3.5 Tempered and regular posteriors
Hereafter, we present a result demonstrating the closeness of the tempered posterior to the regular posterior when $\alpha$ is close to 1.
For two density functions $g_{1}$ and $g_{2}$ with respect to a measure $\mu$, the total variation distance between them is given by
[…]
where $G_{1}$ and $G_{2}$ denote the corresponding probability measures.
Theorem 3.4.
Consider model (2.1) with $\Omega\sim\Pi_{\Omega}.$ Then,
$\lim_{\alpha\to 1_{-}}\|\Pi_{n,\alpha}(\Omega\mid Y)-\Pi_{n}(\Omega\mid Y)\|_{\mathrm{TV}}=0,$
for every $Y\sim P_{\Omega_{0}}$.
It is observed that [7] established a weak convergence result under a more general framework, while Theorem 3.4 offers a significant enhancement by demonstrating strong convergence for the Gaussian likelihood function under consideration. It is also important to note that a similar result to Theorem 3.4 was achieved in Theorem 3 of [13] for a linear Gaussian multiple regression, and our proof is based on their approach. In Bayesian asymptotics, the total variation distance is frequently employed to justify posterior merging of opinions, meaning that the total variation distance between two posterior distributions derived from different priors diminishes as the sample size increases. Theorem 3.4 follows a similar concept, with the distinction that the merging of opinions occurs under small perturbations of the likelihood function.
4 Concluding remarks
In this work, we have investigated the concentration properties of the tempered posterior distribution when estimating precision matrices using the horseshoe prior. Unlike prior studies that employed either a horseshoe-like prior or a simplified horseshoe prior with a fixed shrinkage parameter, our analysis focuses on the fully specified horseshoe prior. We extend the previous theoretical works by considering a high-dimensional setting ($p>n$) and assuming a sparse underlying precision matrix. Furthermore, we also present a novel important aspect by deriving theoretical results for the scenario of model misspecification. We establish a general oracle inequality for the posterior under an abstract setting that encompasses potential model deviations.
Although we have presented comprehensive theoretical results on the concentration of the tempered posterior, several areas remain unexplored. Our findings are primarily derived using the Frobenius norm, so extending these results to other norms represents a valuable avenue for future research. Additionally, exploring results on support recovery, similar to those in [43], would be another promising direction for further investigation.
Acknowledgments
The author acknowledges support from the Centre for Geophysical Forecasting, Norwegian Research Council grant no. 309960, at NTNU.
Conflicts of interest/Competing interests
The author declares no potential conflict of interests.
Appendix A Proofs
A.1 Proof of Theorem 3.1
Proof.
We can check the hypotheses on the KL between the likelihood terms as required in Theorem 2.6 in [2]. For
$K(p_{1},p_{2})=\int p_{1}\log(p_{1}/p_{2}),$
and let $d_{1},\ldots,d_{p}$ denote the eigenvalues of $\Omega_{0}^{-1/2}\Omega\Omega^{-1/2}_{0}$. Then, using Lemma 2, we have,
$K(p_{\Omega_{0}},p_{\Omega})=-\dfrac{1}{2}\sum_{i=1}^{p}\log d_{i}-\dfrac{1}{2}\sum_{i=1}^{n}(1-d_{i}),.$
As $\sum_{i=1}^{n}(1-d_{i})^{2}=\|I_{p}-\Omega_{0}^{-1/2}\Omega\Omega^{-1/2}_{0}\|_{2}^{2},$ when $\|I_{p}-\Omega_{0}^{-1/2}\Omega\Omega^{-1/2}_{0}\|_{2}^{2}$ is small, we have, $\max_{1\leq i\leq p}|1-d_{i}|<1.$ This leads to that,
$K(p_{\Omega_{0}},p_{\Omega})\lesssim\sum_{i=1}^{n}(1-d_{i})^{2}=$ $\|I_{p}-\Omega_{0}^{-1/2}\Omega\Omega^{-1/2}_{0}\|_{2}^{2}$
$=$ $\|\Omega_{0}^{-1/2}(\Omega-\Omega_{0})\Omega^{-1/2}_{0}\|_{2}^{2}$
$\leq$ $\|\Omega_{0}^{-1}\|_{2}^{2}\|\Omega-\Omega_{0}\|_{2}^{2}$
$\leq$ $\varepsilon_{0}^{-2}\|\Omega-\Omega_{0}\|_{2}^{2}.$
Integrating with respect to $\rho_{n}\propto\mathbf{1}_{\|\Omega-\Omega_{0}\|_{2}<\delta}\Pi_{HS}$ where $\delta=[s\log(p/s)/n]^{1/2}$, we have that
$\int K(p_{\Omega_{0}},p_{\Omega})\rho_{n}({\rm d}\Omega)\leq 2\varepsilon_{0}^{-2}\delta^{2}\leq 2\varepsilon_{0}^{-2}\frac{s\log(p/s)}{n}.$ (A.1)
Applying Lemma 1 with $d=p(p-1)/2$, we have that
$\frac{1}{n}\mathcal{K}(\rho_{n},\pi)\leq\frac{1}{n}\log\frac{1}{\Pi_{HS}(\|\Omega-\Omega_{0}\|_{2}<\delta)}\leq\frac{1}{n}Ks\log(p/s).$
Consequently, we can now apply Theorem 2.6 from [2], with
$\varepsilon_{n}=2\varepsilon_{0}^{-2}\frac{s\log(p/s)}{n}.$
The proof is completed.
Proof of Corollary 3.1.
From [39], we have that
$h^{2}(p_{\Omega},p_{\Omega_{0}})\leq D_{1/2}(P_{\Omega},P_{\Omega_{0}})\leq D_{\alpha}(P_{\Omega},P_{\Omega_{0}}),$
for $\alpha\in[0.5,1)$. In addition, from [39], all $\alpha$-Rényi divergences are all equivalent for $10<\alpha<$ 1, through the formula $\frac{\alpha}{\beta}\frac{1-\beta}{1-\alpha}D_{\beta}\leq D_{\alpha}\leq D_{\beta}$ for $\alpha\leq\beta$, thus we have that
$D_{1/2}(P_{\theta},P_{\theta_{0}})\leq\frac{(1-\alpha)1/2}{\alpha(1-1/2)}D_{\alpha}(P_{\theta},P_{\theta_{0}})=\frac{(1-\alpha)}{\alpha}D_{\alpha}(P_{\theta},P_{\theta_{0}}),$
for $\alpha\in(0,0.5)$. Thus, with $c_{\alpha}:=1$ for $\alpha\in[0.5,1)$ and $c_{\alpha}:=(1-\alpha)/\alpha$ for $\alpha\in(0,0.5)$, we obtain from Theorem 3.1 that
$\mathbb{E}\left[\int\frac{1}{c_{\alpha}}h^{2}(p_{\Omega},p_{\Omega_{0}})\pi_{n,\alpha}({\rm d}\Omega)\right]\leq\frac{1+\alpha}{1-\alpha}\varepsilon_{n}.$
From Lemma 3, one gets that
$\mathbb{E}\left[\int\frac{c_{0}}{c_{\alpha}}\|\Omega_{1}-\Omega_{2}\|_{2}^{2}\pi_{n,\alpha}({\rm d}\Omega)\right]\leq\frac{1+\alpha}{1-\alpha}\varepsilon_{n}.$
The proof is completed.
A.2 Proof of Theorem 3.2
Proof.
We utilize the general framework of posterior concentration rates by confirming the criteria outlined in Theorem 2.4 of [2]. This involves assessing the prior concentration rate within Kullback–Leibler neighborhoods.
Put
$K(p_{1},p_{2})=\int p_{1}\log(p_{1}/p_{2}),\text{ and }V(p_{1},p_{2})=\int p_{1}\log^{2}(p_{1}/p_{2}).$
and let $d_{1},\ldots,d_{p}$ denote the eigenvalues of $\Omega_{0}^{-1/2}\Omega\Omega^{-1/2}_{0}$. Controlling the term $K(p_{1},p_{2})=\int p_{1}\log(p_{1}/p_{2})$ is presented in the proof of Theorem 3.1, given in page A.1. Now, using Lemma 2, we have,
$V(p_{\Omega_{0}},p_{\Omega})=\dfrac{1}{2}\sum_{i=1}^{n}(1-d_{i})^{2}+K(p_{\Omega_{0}},p_{\Omega})^{2}.$ (A.2)
As $\sum_{i=1}^{n}(1-d_{i})^{2}=\|I_{p}-\Omega_{0}^{-1/2}\Omega\Omega^{-1/2}_{0}\|_{2}^{2},$ when $\|I_{p}-\Omega_{0}^{-1/2}\Omega\Omega^{-1/2}_{0}\|_{2}^{2}$ is small, we have, $\max_{1\leq i\leq p}|1-d_{i}|<1.$ This gives
$V(p_{\Omega_{0}},p_{\Omega})$ $\lesssim\sum_{i=1}^{n}(1-d_{i})^{2}$
$=$ $\|I_{p}-\Omega_{0}^{-1/2}\Omega\Omega^{-1/2}_{0}\|_{2}^{2}$
$=$ $\|\Omega_{0}^{-1/2}(\Omega-\Omega_{0})\Omega^{-1/2}_{0}\|_{2}^{2}$
$\leq$ $\|\Omega_{0}^{-1}\|_{2}^{2}\|\Omega-\Omega_{0}\|_{2}^{2}$
$\leq$ $\varepsilon_{0}^{-2}\|\Omega-\Omega_{0}\|_{2}^{2}.$
Integrating with respect to $\rho_{n}\propto\mathbf{1}_{\|\Omega-\Omega_{0}\|_{2}<\delta}\Pi_{HS}$ where $\delta=[s\log(p/s)/n]^{1/2}$, we have that
$\int V(p_{\Omega_{0}},p_{\Omega})\rho_{n}({\rm d}\Omega)\leq 2\varepsilon_{0}^{-2}\delta^{2}\leq 2\varepsilon_{0}^{-2}\frac{s\log(p/s)}{n}.$
and from (A.1),
$\int K(p_{\Omega_{0}},p_{\Omega})\rho_{n}({\rm d}\Omega)\leq 2\varepsilon_{0}^{-2}\delta^{2}\leq 2\varepsilon_{0}^{-2}\frac{s\log(p/s)}{n}.$
From Lemma 1, we have that
$\frac{1}{n}\mathcal{K}(\rho_{n},\pi)$ $\leq\frac{1}{n}\log\frac{1}{\Pi_{HS}(\|\Omega-\Omega_{0}\|_{2}<\delta)}$
$\leq\frac{1}{n}Ks\log(p/s).$
Consequently, we can now apply Theorem 2.4 from [2], with
$\varepsilon_{n}=2\varepsilon_{0}^{-2}\frac{s\log(p/s)}{n}.$
The proof is completed. ∎
Proof of Corollary 3.2.
The proof is as in the proof of Corollary 3.1. ∎
Lemma 2 (Lemma B1 in [36]).
Let $p_{k}$ denote the density of a $\mathcal{N}_{d}(0,\Sigma_{k})$ random variable, $k=1,2.$ Denote the corresponding precision matrices by $\Omega_{k}=\Sigma_{k}^{-1},k=1,2.$ Then,
$\mathbb{E}_{p_{1}}\left\{\log\dfrac{p_{1}}{p_{2}}(X)\right\}$ $=\dfrac{1}{2}\left\{\log\det\Omega_{1}-\log\det\Omega_{2}+trace(\Omega^{-1}_{1}\Omega_{2}-I_{d})\right\},$
${\rm Var}_{p_{1}}\left\{\log\dfrac{p_{1}}{p_{2}}(X)\right\}$ $=\dfrac{1}{2}\,trace\{(\Omega_{1}^{-1/2}\Omega_{2}\Omega_{1}^{-1/2}-I_{d})^{2}\}.$
Lemma 3 (Lemma A.1 in [5]).
Let $p_{k}$ denote the density of a $\mathcal{N}_{d}(0,\Sigma_{k})$ random variable, $k=1,2.$ Denote the corresponding precision matrices by $\Omega_{k}=\Sigma_{k}^{-1},k=1,2.$ Then, there exists a positive constant $c_{0}$ such that
$\|\Omega_{1}-\Omega_{2}\|_{2}^{2}\leq c_{0}h^{2}(p_{1},p_{2})$
A.3 Proof of Lemma 1
Proof.
Using the scale-mixture formulation of $\Pi_{HS}$,
$\begin{split}\Pi_{HS}(\|\theta-\theta_{0}\|_{2}<\delta)&=\int_{\tau}pr(\|\theta-\theta_{0}\|_{2}<\delta\mid\tau)f(\tau){\rm d}\tau\\ &\geq\int_{I_{\tau_{*}}}pr(\|\theta-\theta_{0}\|_{2}<\delta\mid\tau)f(\tau){\rm d}\tau,\end{split}$
where $I_{\tau_{*}}=[\tau_{*}/2,\tau_{*}]$ with $\tau_{*}=s(s/p)^{3/2}\{\log(d/s)/n\}^{1/2}$. Let $S=\{1\leq j\leq d:\theta_{0j}\neq 0\}$. For $\tau\in I_{\tau^{*}}$, one can lower bound the conditional probability as
$\begin{split}&pr(\|\theta-\theta_{0}\|_{2}<\delta\mid\tau)\\ &\geq pr\left(\|\theta_{S}-\theta_{0S}\|_{2}<\delta/2\mid\tau\right)pr\left(\|\theta_{S^{c}}\|_{2}<\delta/2\mid\tau\right)\\ &\geq\prod_{j\in S}pr\left(|\theta_{j}-\theta_{0j}|<\frac{\delta}{2s}\mid\tau\right)\prod_{j\in S^{c}}pr\left(|\theta_{j}|<\frac{\delta}{2\sqrt{d}}\mid\tau\right).\end{split}$ (A.3)
For a fixed $\tau\in I_{\tau_{*}}$, we will lower bounds each of the terms in the right hand side of (A.3).
Initially, we examine $pr\{\,|\theta_{j}|<\delta/(2d^{1/2})\mid\tau\}$ with $\tau$ in $I_{\tau_{*}}$. Given $\tau$ and $\lambda$, $\theta_{j}$ follows a normal distribution $N(0,\lambda_{j}^{2}\tau^{2})$. Leveraging the Chernoff bound for a Gaussian random variable, we obtain:
$pr\big{\{}|\theta_{j}|>\delta/(2d^{1/2})\mid\lambda_{j},\tau\big{\}}\leq 2e^{-\delta^{2}/(8p\lambda_{j}^{2}\tau^{2})}\leq 2e^{-\delta^{2}/(8p\lambda_{j}^{2}\tau_{*}^{2})}.$
Thus,
$pr\big{\{}|\theta_{j}|<\delta/(2d^{1/2})\mid\tau\big{\}}$
$=\int_{\lambda_{j}}pr\big{\{}|\theta_{j}|<\delta/(2d^{1/2})\mid\lambda_{j},\tau\big{\}}\,f(\lambda_{j})\,{\rm d}\lambda_{j}$
$\geq\int_{\lambda_{j}}\left\{1-2\exp{\left(-\delta^{2}/(8d\lambda_{j}^{2}\tau_{*}^{2})\right)}\right\}\,f(\lambda_{j})\,{\rm d}\lambda_{j}$
$=1-\dfrac{4}{\pi}\int_{\lambda_{j}}\exp{\left(-\delta^{2}/(8d\lambda_{j}^{2}\tau_{*}^{2})\right)}(1+\lambda_{j}^{2})^{-1}{\rm d}\lambda_{j}=1-\dfrac{4}{\pi}\,\mathcal{I}.$
We now upper-bound the integrand as follows,
$\mathcal{I}$ $=\int_{\lambda_{j}}\exp{\left(-\dfrac{\delta^{2}}{8d\tau_{*}^{2}\lambda_{j}^{2}}\right)}\,(1+\lambda_{j}^{2})^{-1}{\rm d}\lambda_{j}$
$\leq\int_{\lambda_{j}}\exp{\left(-\dfrac{\delta^{2}}{8d\tau_{*}^{2}\lambda_{j}^{2}}\right)}\lambda_{j}^{-2}{\rm d}\lambda_{j}$
$=\frac{1}{2}\int_{0}^{\infty}z^{-1/2}\exp{\left(-\dfrac{\delta^{2}z}{8d\tau_{*}^{2}}\right)}{\rm d}z,$
$=\dfrac{\Gamma(1/2)}{\{2\delta^{2}/(8d\tau_{*}^{2})\}^{1/2}}=\dfrac{2s^{2}\sqrt{\pi}}{d},$
where a substitution $z=1/\lambda^{2}$ at the third step was made. Thus, for $\tau\in I_{\tau_{*}}$,
$pr(|\theta_{j}|<\delta/2d^{1/2}\mid\tau)\geq 1-\dfrac{8s^{2}}{d\sqrt{\pi}}.$
Now, we lower bound $pr(|\theta_{j}-\theta_{0j}|<\delta_{0}|\tau)$ with $\tau\in I_{\tau_{*}}$. Letting $\delta_{0}=\delta/(2\sqrt{s})$, then
$\begin{split}&pr(|\theta_{j}-\theta_{0j}|<\delta_{0}\mid\tau)\\ &=(\frac{2}{\pi^{3}})^{1/2}\int_{\lambda_{j}}\int_{\mid\theta_{j}-\theta_{0}\mid<\delta_{0}}\exp\{{-\theta_{j}^{2}/(2\lambda_{j}^{2}\tau^{2})}\}\dfrac{1}{\lambda_{j}\tau(1+\lambda_{j}^{2})}\,{\rm d}\theta_{j}{\rm d}\lambda_{j}\\ &\geq(\frac{2}{\pi^{3}})^{1/2}\int_{|\theta_{j}-\theta_{0}|<\delta_{0}}\int_{1/\tau}^{2/\tau}\exp\{{-\theta_{j}^{2}/(2\lambda_{j}^{2}\tau^{2})}\}\dfrac{1}{\lambda_{j}\tau(1+\lambda_{j}^{2})}\,{\rm d}\lambda_{j}{\rm d}\theta_{j}\\ &\geq(\frac{2}{\pi^{3}})^{1/2}\int_{|\theta_{j}-\theta_{0}|<\delta_{0}}\exp(-\theta_{j}^{2}/2)\left(\int_{1/\tau}^{2/\tau}\dfrac{1}{1+\lambda_{j}^{2}}\,{\rm d}\lambda_{j}\right){\rm d}\theta_{j},\\ \end{split}$
since for $\lambda_{j}\in[1/\tau,2/\tau]$, $21/(\lambda_{j}\tau)\geq 1/$ 2 and $\exp\{{-\theta_{j}^{2}/(2\lambda_{j}^{2}\tau^{2})}\}\geq\exp{(-\theta_{j}^{2}/2)}$. Continuing,
$pr(|\theta_{j}-\theta_{0j}|<\delta_{0}\mid\tau)$ $\geq(2/\pi^{3})^{1/2}\,\dfrac{\tau}{4+\tau^{2}}\,\int_{|\theta_{j}-\theta_{0}|<\delta_{0}}\exp{(-\theta_{j}^{2}/2)}{\rm d}\theta_{j}$
$\geq(2/\pi^{3})^{1/2}\,\,\dfrac{\tau}{4+\tau^{2}}\,\,\exp\{{-(M+1)^{2}/2}\}\,\,\delta_{0}$
$\geq K\,\tau\,\delta_{0}\geq K\,\frac{\tau_{*}\delta}{4\sqrt{s}}\geq Ks\left(\frac{s}{p}\right)^{3/2}\frac{\log(p/s)}{n}$
$\geq K\left(\frac{s}{p}\right)^{5/2}=Ke^{-(5/2)\log(p/s)},$
where in the third step, we used $54+\tau^{2}<$ 5 and $n<p$ in the final step.
By substituting these bounds into (A.3), we obtain, for $\tau\in I_{\tau_{*}}$,
$pr(\|\theta-\theta_{0}\|_{2}<\delta\mid\tau)$
$\geq\left(Ke^{-(5/2)\log(p/s)}\right)^{s}\left(1-\dfrac{8s^{2}}{p\sqrt{\pi}}\right)^{p-s}$
$\geq e^{-Ks\log(p/s)},$
where $K$ is a positive constant. The proof concludes by noting that the probability of $\tau$ being within the interval $I_{\tau_{*}}$ is greater than or equal to $\tau_{*}/(2\pi)$. Thus, with a minor deviation in notation, we obtain,
$\Pi_{HS}(\|\theta-\theta_{0}\|_{2}<\delta)\geq e^{-Ks\log(p/s)},$
for some positive constant $K$. The proof is completed. ∎
A.4 Proof of Theorem 3.3
Proof.
We can check the assumptions as in Theorem 2.7 in [2].
Let us define $A=\Omega_{1}^{-1/2}\Omega_{2}\Omega_{1}^{-1/2}.$ Note that, for a random variable $Z\sim\mathcal{N}_{d}(0,\Sigma)$, we have, $\mathbb{E}(Z^{T}AZ)=trace(A\Sigma).$ Then, for $X\sim\mathcal{N}_{d}(0,\Sigma_{0})$,
$\mathbb{E}_{p_{0}}\left\{\log\dfrac{p_{1}}{p_{2}}(X)\right\}$ $=\dfrac{1}{2}\left\{\log\det\Omega_{1}-\log\det\Omega_{2}+\mathbb{E}_{p_{0}}(X^{T}(\Omega_{2}-\Omega_{1})X)\right\}$
$=\dfrac{1}{2}\left\{\log\det\Omega_{1}-\log\det\Omega_{2}+trace[(\Omega_{2}-\Omega_{1})\Sigma_{0}]\right\}.$
Thus, we have that
$\mathbb{E}_{\Omega_{0}}\left[\log\frac{p_{\Omega_{*}}}{p_{\Omega}}(X)\right]$ $\leq\dfrac{1}{2}\left\{\log\left(\frac{\det(\Omega_{*})}{\det(\Omega)}\right)+trace[(\Omega-\Omega_{*})\Omega_{0}^{-1}]\right\}$
$\leq\dfrac{1}{2}\left\{\log\left(\frac{\det(\Omega_{*})}{\det(\Omega)}\right)+\|(\Omega-\Omega_{*})\Omega_{*}^{-1}\Omega_{*}\|_{2}\|\Omega_{0}^{-1}\|_{2}\right\}$
$\leq\dfrac{1}{2}\left\{\log\left(\frac{\det(\Omega_{*})}{\det(\Omega)}\right)+\|\Omega\Omega_{*}^{-1}-I_{p}\|_{2}\|\Omega_{*}\|_{2}\varepsilon_{0}^{-1}\right\}.$
Let $d_{1},\ldots,d_{p}$ denote the eigenvalues of $\Omega\Omega_{*}^{-1}$. As $\sum_{i=1}^{n}(1-d_{i})^{2}=\|I_{p}-\Omega\Omega_{*}^{-1}\|_{2}^{2},$ when $\|I_{p}-\Omega\Omega_{*}^{-1}\|_{2}^{2}$ is small, we have, $\max_{1\leq i\leq p}|1-d_{i}|<1.$ This leads to that,
$\mathbb{E}_{\Omega_{0}}\left[\log\frac{p_{\Omega_{*}}}{p_{\Omega}}(X)\right]$ $\leq\dfrac{1}{2}\left\{\sum_{i=1}^{p}\log d_{i}+\|\Omega_{*}\|_{2}\varepsilon_{0}^{-1}\sum_{i=1}^{p}(1-d_{i})\right\}$
$\leq K\|\Omega_{*}\|_{2}\varepsilon_{0}^{-1}\sum_{i=1}^{p}(1-d_{i})^{2}$
$\leq K_{\varepsilon_{0}}\|\Omega-\Omega_{*}\|_{2}^{2},$
for some absolute constant $K_{\varepsilon_{0}}>0$ depending only on $\varepsilon_{0}$. Integrating with respect to $\rho_{n}\propto\mathbf{1}_{\|\Omega-\Omega_{*}\|_{2}<\delta}\Pi_{HS}$ where $\delta=[s\log(p/s)/n]^{1/2}$, we have that
$\int\mathbb{E}_{\Omega_{0}}\left[\log\frac{p_{\Omega_{*}}}{p_{\Omega}}(X)\right]\rho_{n}({\rm d}\Omega)\leq K_{\varepsilon_{0}}\delta^{2}\leq K_{\varepsilon_{0}}\frac{s\log(p/s)}{n}.$
To apply Theorem 2.7 in [2] it remains to compute the KL between the approximation $\rho_{n}$ and the prior. From Lemma 1, we have that
$\frac{1}{n}\mathcal{K}(\rho_{n},\pi)$ $\leq\frac{1}{n}\log\frac{1}{\Pi_{HS}(\|\Omega-\Omega_{0}\|_{2}<\delta)}$
$\leq\frac{1}{n}Ks\log(p/s).$
To obtain an estimate of the rate $\varepsilon_{n}$ as in Theorem 2.7 in [2], we put together those bounds and choosing
$\varepsilon_{n}=K_{\varepsilon_{0}}\frac{s\log(p/s)}{n}.$
The proof is completed.
A.5 Proof of Theorem 3.4
Proof.
The proof follows an argument as in the proof of Theorem 3 of [13].
Put
$m_{\alpha}(Y)=\int L(Y|\Omega)^{\alpha}\,\Pi_{\Omega}(d\Omega)$
where $\alpha\in(0,1)$ and
$m(Y)=\int L(Y|\Omega)\,\Pi_{\Omega}(d\Omega).$
One has that $m(Y)>0$ for every $Y$ due to the positivity of $L(Y|\Omega)$ and $\Pi_{\Omega}(d\Omega)$. For every $\Omega$,
$\lim_{\alpha\to 1_{-}}L(Y|\Omega)^{\alpha}\,\Pi_{\Omega}(d\Omega)=L(Y|\Omega)\,\Pi_{\Omega}(d\Omega).$
Since
$L(Y|\Omega)^{\alpha}\,\Pi_{\Omega}(d\Omega)\leq\Pi_{\Omega}(d\Omega)$
and $\int\Pi_{C}(dC)\Pi_{\Sigma}(d\Sigma)=1$, by applying the dominated convergence theorem, it follows that
$\lim_{\alpha\to 1_{-}}m_{\alpha}(Y)=m(Y).$
Combining, we get that $\lim_{\alpha\to 1_{-}}\Pi_{n,\alpha}(\Omega\mid Y)=\Pi_{n}(\Omega\mid Y)$ for all $\Omega$. Then, by applying Scheffé’s theorem [24, 37], we obtain the desired result. ∎